OpenStudy (anonymous):

Simplify the given expression. square root of negative 10 end root times square root of negative 8

OpenStudy (anonymous):

negative 4 square root 5 4 square root of 5 4 I square root of 5 negative 4 I square root of 5

OpenStudy (anonymous):

i got the 3rd one is it right?

OpenStudy (anonymous):

i∗10−−√∗8√∗i −80−−√=−45√

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{-10}*\sqrt{-8}\] \[\large \sqrt{-1*10}*\sqrt{-1*8}\] \[\large \sqrt{-1}*\sqrt{10}*\sqrt{-1}*\sqrt{8}\] \[\large \sqrt{-1}*\sqrt{-1}*\sqrt{10}*\sqrt{8}\] \[\large i*i*\sqrt{10}*\sqrt{8}\] \[\large i^2*\sqrt{10}*\sqrt{8}\] \[\large -1*\sqrt{10}*\sqrt{8}\] I'll let you finish up.

OpenStudy (jdoe0001):

\(\bf \sqrt{-10}\cdot \sqrt{-8}\implies \sqrt{10\cdot -1}\cdot \sqrt{8\cdot -1}\implies \sqrt{10}\cdot \sqrt{-1}\cdot \sqrt{8}\cdot \sqrt{-1} \\ \quad \\ \sqrt{10}\ i\cdot \sqrt{8}\ i\implies \sqrt{10\cdot 8}\ i \cdot i\implies \sqrt{80}\ i^2\implies \sqrt{4^2\cdot 5}\ i^2 \\ \quad \\ {\color{brown}{ i^2\to \sqrt{-1}\cdot \sqrt{-1}\to \sqrt{(-1)^2}\to -1}} \\ \quad \\ -1\cdot \sqrt{4^2\cdot 5}\implies -4\sqrt{5}\)